# EP-Diffuser: An Efficient Diffusion Model for Traffic Scene Generation and Prediction via Polynomial Representations

### v1

@@ Abstract
As the prediction horizon increases, predicting the future evolution of traffic scenes becomes increasingly difficult due to the multi-modal nature of agent motion. Most state-of-the-art (SotA) prediction models primarily focus on forecasting the most likely future. However, for the safe operation of autonomous vehicles, it is equally important to cover the distribution for plausible motion alternatives. To address this, we introduce EP-Diffuser, a novel parameter-efficient diffusion-based generative model designed to capture the distribution of possible traffic scene evolutions. Conditioned on road layout and agent history, our model acts as a predictor and generates diverse, plausible scene continuations. We benchmark EP-Diffuser against two SotA models in terms of accuracy and plausibility of predictions on the Argoverse 2 dataset. Despite its significantly smaller model size, our approach achieves both highly accurate and plausible traffic scene predictions. We further evaluate model generalization ability in an out-of-distribution (OoD) test setting using Waymo Open dataset and show superior robustness of our approach. The code and model checkpoints can be found here: https://github.com/continental/EP-Diffuser.
I Introduction
@@
Public motion datasets, such as Argoverse 2 (A2) [1] and Waymo Open (WO) [2], provide real-world traffic scene data and host associated motion prediction competitions to advance research in this field. The evolution of traffic scenes over long time horizons is governed by an inherently multi-modal probability distribution, as multiple plausible futures exist depending on road topology and agent interactions. However, motion datasets can only record a single observed future sample (the ground truth) per scene. As a result, motion prediction competitions typically frame the problem as a regression problem, where models are trained to estimate the most likely outcome based on available ground truth data. Existing approaches can be categorized into two main types:
@@
Regression-based metrics, such as Average Displacement Error (ADE), Final Displacement Error (FDE), and their variants, are widely used in these competitions. These metrics evaluate prediction accuracy by measuring how closely a prediction matches a single ground truth trajectory. However, this evaluation approach presents two key limitations illustrated in Figure 1: First, it does not measure the diversity and coverage of predictions. Second, it fails to account for plausibility and consistency, both of which are essential for safe trajectory planning. Furthermore, focusing solely on the most probable future evolution while ignoring other plausible possibilities may induce overconfident and risky behavior, making it insufficient for planning algorithms, as highlighted in recent trajectory planning studies [3, 4, 5].
To address these challenges, recent works have begun exploring more comprehensive frameworks using generative models. These generative approaches aim to learn the distribution of future traffic scenes under given conditions (i.e., observed agent history and road layout) and enable sampling from the modeled joint distribution – a task referred to as traffic scene generation. Therefore, these approaches can capture not only the most likely traffic scene evolution but also a range of plausible alternatives.
@@
Another fundamental challenge is assessing model generalization. Prior studies typically evaluate model performance using a test split from the same dataset used for training. While motion datasets attempt to ensure disjoint splits between training and testing, these subsets still share underlying biases, such as recurring road layouts, traffic flow patterns, and artifacts introduced during data collection and pre-processing. Generative models are known to exhibit memorization – the tendency to produce near-replicas of training data [9, 10]. Consequently, models tested solely on these similar samples may appear to perform well by leveraging their memorizing capability instead of learning robust, transferable representations of traffic patterns.
@@
With these considerations in mind, we present EP-Diffuser, a generative model for traffic scene generation conditioned on road layout and observed agent history, thereby performing joint prediction. Unlike traditional models that incorporate sequence-based data, such as lists of trajectory observations or map points, our model employs polynomial representations for both map elements and trajectories on the model’s input and output sides. To comprehensively evaluate our approach, we extend beyond regression-based metrics and benchmark EP-Diffuser against two state-of-the-art (SotA) models from a plausibility perspective. Specifically, we incorporate Waymo’s “Sim Agents” metrics [13] and assess performance under OoD scenarios from the WO dataset. Our results demonstrate that the polynomial representation enhances the efficiency of the denoising process, temporal consistency in generated agent kinematics, and generalization in OoD test settings. Our contributions are summarized as follows:
@@
Our paper is organized as follows: We first review recent traffic scene prediction and generation models, along with their evaluation metrics. We then introduce two benchmark models and the “Sim Agents” metrics, highlighting key differences from regression-based metrics. Next, we present our diffusion-based approach with constrained parametric representations. We evaluate our model against benchmarks, analyzing both plausibility and regression-based performance. Finally, we extend our evaluation to OoD scenes to assess generalization beyond the training data.
@@ II-A Traffic Scene Prediction and Benchmark Models
To evaluate OoD generalization, we focus on models trained on the smaller A2 dataset, aiming for a more rigorous and insightful assessment of their robustness when applied to scenes from the larger WO dataset. While many open-source models exist for the multi-modal marginal prediction task, there are far fewer open-source benchmark models specifically addressing multi-modal traffic scene (joint) prediction with documented performance and reproducible results. This limitation narrows the pool of suitable benchmark models for our experiment.
@@ II-B Sim Agents Metrics
In contrast to regression-based tasks, “Sim Agents” frames traffic scene prediction as a multi-agent generative task, emphasizing the importance of capturing the diversity and realism of traffic behaviors [13]. Rather than solely minimizing prediction errors, models are evaluated on their ability to produce realistic and socially consistent traffic scenes.
To assess these qualities, the challenge employs a comprehensive evaluation framework comprising the following metrics:
@@ III Data Representation and Model
We introduce Everything Polynomial Diffuser (EP-Diffuser), where polynomial inputs and outputs serve as the key innovation in our diffusion model. Unlike other diffusion models that use sequence-based representations [7, 18, 6], our approach integrates polynomial representations for map elements and trajectories. This significantly simplifies the diffusion-denoising process by reducing data dimensionality. Moreover, polynomial representations effectively regularize measurement noise and inherently ensure temporal consistency in predicted trajectories, addressing a key challenge in current SotA models.
In the subsequent sections, we first describe our approach to represent diverse data types using polynomial representations, followed by the implementation details of our model.
@@ III-A Data as Polynomials
We employ Bernstein polynomials to represent the agent histories, future trajectories, and map geometry. The parameters of Bernstein polynomials have spatial semantics as control points. The A2 dataset segments each 11-second recording into a 5-second history and a 6-second future. We represent different data types as follows:
Agent History: Following the Akaike Information Criterion (AIC) [22] from the study [23], we represent 5-second history trajectories of vehicles, cyclists, and pedestrians in A2 using optimal 5-degree polynomials. We also use the 5-degree polynomial for the ego vehicle. We track the control points of agent history with the method proposed in [24] and incorporate the observation noise models proposed in [23].
@@ III-B Model Architecture
The entire pipeline of EP-Diffuser is illustrated in Figure 2. EP-Diffuser employs an encoder-denoiser architecture similar to [8] but with different data representations. Implementation details are provided in Appendix A.
Diffusion models typically require many denoising steps to reconstruct outputs from sampled noise [27]. To accelerate this process, we adopt the Denoising Diffusion Implicit Models (DDIM) method [20], which reduces the number of denoising steps without compromising sample quality (see Appendix A-A for details). Consistent with the configuration of OptTrajDiff, we also use 10 denoising steps for EP-Diffuser.
@@ IV-A Experiment Setup
IV-A1 Training and Testing on A2
Following the Argoverse 2 Motion Prediction Competition setup, models are required to output 6-second predictions given 5-second history. All models are trained with this setup from scratch with their original hyperparameters on the A2 training split containing scenarios.
@@ IV-A1 Training and Testing on A2
Due to the computational cost of “Sim Agents” metric calculations, we focus on the 500 most challenging traffic scenes in A2 validation set. We identify these scenes based on the largest deviations between the ground truth and the constant velocity model, as measured by the realism meta metric.
train / test | model | realism | kinematic | interactive | map-based | minADE | # model params |
meta metric | metrics | metrics | metrics | [m] | [million] | ||
ground truth | 0.841 | 0.619 | 0.851 | 0.956 | 0 | - | |
constant velocity | 0.194 | 0.261 | 0.175 | 0.180 | 2.713 | - | |
A2 / A2 | FMAE-MA [14] (ensemble) | 0.636 | 0.320 | 0.679 | 0.760 | 0.479 | 11.4 (61.9) |
OptTrajDiff [7] | 0.709 | 0.459 | 0.717 | 0.841 | 0.449 | 12.5 | |
EP-Diffuser (ours) | 0.713 | 0.507 | 0.707 | 0.838 | 0.546 | 3.0 | |
ground truth | 0.837 | 0.585 | 0.866 | 0.944 | 0 | - | |
constant velocity | 0.175 | 0.162 | 0.176 | 0.180 | 1.498 | - | |
A2 / WO | FMAE-MA [14] (ensemble) | 0.630 | 0.271 | 0.716 | 0.723 | 0.426 | 11.4 (61.9) |
OptTrajDiff [7] | 0.721 | 0.403 | 0.771 | 0.839 | 0.357 | 12.5 | |
EP-Diffuser (ours) | 0.742 | 0.456 | 0.782 | 0.854 | 0.372 | 3.0 |
IV-A2 OoD Testing on WO
@@
We apply the same sampling strategies and test models on the 500 most challenging scenes from the WO validation split, using the same selection criteria as in A2 for the homogenized 4.1-second prediction task.
For OoD testing, all three models are trained on the homogenized A2 training split and tested on OoD samples from the homogenized WO validation split. Results are reported based on the 4.1-second prediction.
@@ IV-B Comparison with SotA Models on A2
Table II (upper section) summarizes the results of all three models tested on A2. The ground truth and constant velocity baselines serve as the upper and lower performance bounds, respectively. Figure 3 visualizes multiple samples generated by EP-Diffuser.
We observe a complete reversal in rankings when evaluating prediction plausibility versus . Despite its smaller model size ( million parameters) and the highest (), EP-Diffuser achieves the highest realism meta score at and performs comparably to OptTrajDiff in terms of agent interaction and map adherence. Notably, EP-Diffuser demonstrates superior agent kinematics (), outperforming OptTrajDiff () and FMAE-MA ().
As an illustrative example, Figure 4 visualizes the vehicle kinematics during a left-turn maneuver in A2. EP-Diffuser produces the most plausible agent kinematics according to the measures in [28]. Since polynomials remain polynomials under differentiation, they inherently ensure the temporal consistency of agent kinematics by design.
Figure 5 reports the impact of the number of DDIM denoising steps on realism meta scores and inference time in diffusion-based models, with FMAE-MA as the regression-based baseline for comparison. Both generative models achieve their peak realism scores with relatively few DDIM denoising steps – EP-Diffuser at 5 steps and OptTrajDiff at 10 steps. Increasing the number of denoising steps beyond these points does not improve prediction realism but significantly increases computational cost.
Across all denoising step configurations, EP-Diffuser consistently outperforms OptTrajDiff in both metrics, achieving higher realism scores with lower inference time. While EP-Diffuser’s inference time exceeds FMAE-MA’s, it reaches peak performance with 5 denoising steps in just , making it viable for real-time applications.
@@ IV-C OoD Testing on WO
In the OoD Setting, the model trained on A2 is asked to generate traffic scene continuations for independent scenes taken from the WO dataset – thus removing any sort of shared bias between training and test data.
The OoD testing results are presented in the lower section in Table II. For reference, we also include the ground truth results as the upper bound and the constant velocity baseline as the lower bound.
In the OoD setting, EP-Diffuser maintains the top scores in realism meta and agent kinematics. Additionally, it also demonstrates improved performance by achieving the best agent interactions and map adherence scores (with all subscores detailed in Table III in the Appendix). Furthermore, EP-Diffuser outperforms FMAE-MA in and closely matches OptTrajDiff ( vs. , respectively). These results demonstrate EP-Diffuser’s ability to learn robust, transferable representations from training data, highlighting its enhanced generalization beyond dataset-specific patterns.

### v3

@@ Abstract
As the prediction horizon increases, predicting the future evolution of traffic scenes becomes increasingly difficult due to the multi-modal nature of agent motion. Most state-of-the-art (SotA) prediction models primarily focus on forecasting the most likely future. However, for the safe operation of autonomous vehicles, it is equally important to cover the distribution for plausible motion alternatives. To address this, we introduce EP-Diffuser, a novel parameter-efficient diffusion-based generative model designed to capture the distribution of possible traffic scene evolutions. Conditioned on road layout and agent history, our model acts as a predictor and generates diverse, plausible scene continuations. We benchmark EP-Diffuser against two SotA models in terms of plausibility, diversity, and accuracy of predictions on the Argoverse 2 dataset. Despite its significantly smaller model size, our approach achieves both highly plausible and diverse traffic scene predictions with comparable accuracy. We further evaluate model generalization in an out-of-distribution (OoD) test setting using Waymo Open dataset and show superior robustness of our approach. The code and model checkpoints are available at: https://github.com/continental/EP-Diffuser.
Index Terms:
Autonomous Agents; Deep Learning Methods; Performance Evaluation and BenchmarkingI Introduction
@@
Public motion datasets, such as Argoverse 2 (A2) [1] and Waymo Open (WO) [2], provide real-world traffic scene data and host associated motion prediction competitions to advance research in this field. The evolution of traffic scenes over long time horizons is governed by an inherently multi-modal probability distribution. Multiple plausible futures exist depending on road topology and agent interactions. However, motion datasets can only record a single observed future sample (the ground truth) per scene. As a result, motion prediction competitions typically frame the problem as a regression problem, where models are trained to estimate the most likely outcome based on available ground truth data. Existing approaches can be categorized into two main types:
@@
Regression-based metrics, such as Average Displacement Error (ADE), Final Displacement Error (FDE), and their variants, are widely used in these competitions. These metrics evaluate prediction accuracy by measuring how closely predictions match ground truth trajectories. However, this evaluation approach presents two key limitations illustrated in Figure 1: First, it does not measure the diversity and coverage of predictions. Second, it fails to account for plausibility and consistency, both of which are essential for safe trajectory planning. Furthermore, focusing solely on the most probable future evolution while ignoring other plausible possibilities may induce overconfident and risky behavior, making it insufficient for planning algorithms, as highlighted in recent trajectory planning studies [3, 4, 5, 6].
Regression-based approaches have attempted to improve prediction diversity using discrete trajectory sets [7] or modified training objectives [8, 9]. In parallel, other studies have explored generative modeling frameworks. These include methods based on GANs [10, 11], VAEs [12], and diversity-enhancing sampling methods [13, 14], which initially aim to produce diverse marginal predictions for individual agents.
Building on this trend, recent generative approaches have begun addressing the more challenging task of multi-agent traffic scene generation. These models aim to learn the distribution of future traffic scenes under given conditions (i.e., observed agent history and road layout) and enable sampling from the modeled joint distribution. Therefore, they can capture not only the most likely traffic scene evolution but also a range of plausible alternatives.
@@
Another fundamental challenge is assessing model generalization. Prior studies typically evaluate model performance using a test split from the same dataset used for training. While motion datasets attempt to ensure disjoint splits between training and testing, these subsets still share underlying biases, such as recurring road layouts, traffic flow patterns, and artifacts introduced during data collection and pre-processing. Generative models are known to exhibit memorization – the ability to produce near-replicas of training data [18, 19]. Consequently, models tested solely on these similar samples may appear to perform well by leveraging their memorizing capability instead of learning robust, transferable representations of traffic patterns.
@@
With these considerations in mind, we present Everything Polynomial Diffuser (EP-Diffuser), a generative model for traffic scene generation conditioned on road layout and observed agent history, thereby performing joint prediction. Unlike traditional models that incorporate sequence-based data – such as lists of trajectory observations or map points – or those that use polynomials solely as outputs [22], our model employs polynomial representations for both map elements and trajectories on the model’s input and output sides. To comprehensively evaluate our approach, we extend beyond regression-based metrics and benchmark EP-Diffuser against two state-of-the-art (SotA) models from the perspectives of plausibility, coverage, and accuracy. Specifically, we incorporate Waymo’s “Sim Agents” metrics [23] for evaluating plausibility and assess performance under OoD scenarios from the WO dataset. Our results demonstrate that the polynomial representation enhances the efficiency of the denoising process, temporal consistency in generated agent kinematics, and generalization in OoD test settings. Our contributions are summarized as follows:
@@
Our paper is organized as follows: We first review recent traffic scene prediction and generation models, along with their evaluation metrics. We then introduce two benchmark models and the “Sim Agents” metrics, highlighting key differences from regression-based metrics. Next, we present our diffusion-based approach with constrained parametric representations. We evaluate our model against benchmarks, analyzing plausibility, coverage, and regression-based performance. Finally, we extend our evaluation to OoD scenes to assess generalization beyond the training data.
@@ II-A Traffic Scene Prediction and Benchmark Models
Cross-dataset testing introduces a notable distribution shift [21, 20] and is well suited for evaluating OoD generalization. We focus on models trained on the smaller A2 dataset, enabling a more rigorous assessment of their robustness when applied to scenes from the larger WO dataset. While many open-source models exist for the multi-modal marginal prediction task, there are far fewer open-source benchmark models specifically addressing multi-modal traffic scene (joint) prediction with documented performance and reproducible results. This limitation narrows the pool of suitable benchmark models for our experiment.
@@ II-B Sim Agents Metrics
In contrast to regression-based tasks, “Sim Agents” frames traffic scene prediction as a multi-agent generative task, emphasizing the importance of capturing the diversity and realism of traffic behaviors [23]. Rather than focusing solely on minimizing prediction errors, Sim Agents metrics evaluate models based on their ability to produce plausible and consistent traffic scenes using several complementary metrics:
@@ III Data Representation and Model
EP-Diffuser extends our previous marginal prediction model, EP [20], by introducing a generative diffusion framework and enabling multi-agent joint trajectory generation. Unlike other diffusion models that use sequence-based representations [16, 28, 15], our approach integrates polynomial representations for map elements and trajectories.
This polynomial design offers three key benefits. First, it provides a low-dimensional, truncated Taylor-like approximation that captures key motion patterns without modeling complex high-order residuals. This improves the efficiency of the diffusion-denoising process and simplifies the overall learning task. Second, since polynomial basis functions possess well-defined derivatives, training with positional loss implicitly influences kinematic terms such as velocity and acceleration. This yields smooth and physically plausible trajectories when appropriately regularized with moderate polynomial degrees. Third, it enhances model generalization by mitigating dataset-specific bias and noise commonly present in sequence-based representations, as highlighted in [20]. Unlike data-dependent dimensionality reduction methods such as Principal Component Analysis (PCA) used in [17], polynomial basis functions are data-independent and support better generalization across datasets.
In the subsequent sections, we first describe our approach for representing diverse data types using polynomial representations, followed by the implementation details of our model.
@@ III-A Data as Polynomials
Different types of polynomials of the same degree – such as monomial, Bernstein, and Chebyshev polynomials – can be linearly transformed into one another [32], offering the same expressiveness without affecting the model’s capacity. In this work, we employ Bernstein polynomials to represent agent histories, future trajectories, and map geometry, though our approach is not restricted to this basis. The parameters of Bernstein polynomials have spatial semantics as control points, making them particularly convenient for coordinate transformations such as translation, rotation, and scaling.
The A2 dataset segments each 11-second recording into a 5-second history and a 6-second future. The choice of polynomial degrees for agent trajectories has been extensively analyzed in our prior work [33], which examines the trade-off between representation fidelity and polynomial degree across various agent types and trajectory lengths in both the A2 and WO datasets. Based on the suggestions in [33], we represent different data types as follows:
Agent History: Following the Akaike Information Criterion (AIC) [34] from the study [33], we represent 5-second history trajectories of vehicles, cyclists, and pedestrians in A2 using optimal 5-degree polynomials. We also use the 5-degree polynomial for the ego vehicle. We track the control points of agent history with the method proposed in [32] and incorporate the observation noise models in [33].
@@ III-B Model Architecture
The pipeline of EP-Diffuser is illustrated in Figure 2. The model employs an encoder-denoiser architecture, with implementation details provided in Appendix A. To mitigate variability in sampling procedures and accelerate the inference process, we adopt the Denoising Diffusion Implicit Models (DDIM) method [30] with 10 denoising steps – consistent with the configuration of OptTrajDiff (details in Appendix A-A).
@@ IV-A Experiment Setup
IV-A1 In-Distribution Training and Testing on A2
Following the Argoverse 2 Motion Prediction Competition setup, models are required to output 6-second predictions given 5-second history. All models are trained with this setup from scratch with their original hyperparameters on the A2 training split containing 199,908 scenarios.
| train / test | model | realism | kinematic | interactive | map | minADE | coverage | # model params | ||
| meta | metrics | metrics | metrics | [m] | [m] | [million] | ||||
R20P |
A2 / A2 | FMAE-MA [24] | 0.749 | 0.459 | 0.786 | 0.867 | 0.333 | 2.000 | 11.4 (61.9) | |
| OptTrajDiff [16] | 0.795 | 0.564 | 0.805 | 0.915 | 0.325 | 4.447 | 12.5 | |||
| Seq-Diffuser (ablation) | 0.748 | 0.414 | 0.789 | 0.885 | 0.331 | 1.718 | 3.1 | |||
| EP-Diffuser (ours) | 0.809 | 0.632 | 0.808 | 0.913 | 0.398 | 4.178 | 3.0 | |||
| A2 / WO | FMAE-MA [24] | 0.709 | 0.309 | 0.800 | 0.819 | 0.363 | 3.560 | 11.4 (61.9) | ||
| OptTrajDiff [16] | 0.768 | 0.420 | 0.827 | 0.890 | 0.310 | 3.496 | 12.5 | |||
| Seq-Diffuser (ablation) | 0.688 | 0.212 | 0.801 | 0.814 | 0.458 | 2.446 | 3.1 | |||
| EP-Diffuser (ours) | 0.788 | 0.491 | 0.834 | 0.900 | 0.348 | 3.224 | 3.0 | |||
C500 |
A2 / A2 | FMAE-MA [24] | 0.636 | 0.320 | 0.679 | 0.760 | 0.479 | 2.780 | 11.4 (61.9) | |
| OptTrajDiff [16] | 0.709 | 0.459 | 0.717 | 0.841 | 0.449 | 5.361 | 12.5 | |||
| Seq-Diffuser (ablation) | 0.634 | 0.279 | 0.674 | 0.786 | 0.467 | 2.020 | 3.1 | |||
| EP-Diffuser (ours) | 0.713 | 0.507 | 0.707 | 0.838 | 0.546 | 4.500 | 3.0 | |||
| A2 / WO | FMAE-MA [24] | 0.630 | 0.271 | 0.716 | 0.723 | 0.426 | 3.855 | 11.4 (61.9) | ||
| OptTrajDiff [16] | 0.721 | 0.403 | 0.771 | 0.839 | 0.357 | 3.189 | 12.5 | |||
| Seq-Diffuser (ablation) | 0.630 | 0.210 | 0.739 | 0.730 | 0.448 | 1.980 | 3.1 | |||
| EP-Diffuser (ours) | 0.742 | 0.456 | 0.782 | 0.854 | 0.372 | 2.786 | 3.0 | |||
@@ IV-A1 In-Distribution Training and Testing on A2
IV-A2 Out-of-Distribution Testing on WO
@@
For OoD testing, we apply the same sampling strategies as in A2. All three models are trained on the homogenized A2 training split and tested on OoD samples from the homogenized WO validation split. Results are reported based on the 4.1-second prediction.
IV-A3 Test Data Selection
Due to the computational cost of “Sim Agents” metric calculations, we subsample the validation data from both the A2 and WO datasets under two distinct settings:
20% Random Subsample: We randomly select of the validation set from each dataset, corresponding to 5,000 samples from A2 and 8,400 samples from WO. This setting aims to reflect the model’s overall performance across typical scenes presented by datasets. We refer to this subset as R20P in the remainder of the paper.
500 Most Challenging Scenes: We identify and select the 500 most difficult traffic scenes in each validation set based on the largest deviations between ground truth trajectories and those generated by a constant velocity model, as measured by the realism meta metric. This setting emphasizes model performance under more complex and demanding conditions. We refer to this subset as C500 in the remainder of the paper.
IV-A4 Coverage Metric
Inspired by the inter-policy diversity metric in [37], we compute the average pairwise distance between the final positions of the 32 sampled trajectories to measure the trajectory diversity. This metric serves as an auxiliary indicator and is not used as a standalone measure of performance – since high coverage is only meaningful when accompanied by high plausibility.
@@ IV-B Comparison with SotA Models on A2
Table II (“A2/A2” sections) summarizes the in-distribution results of all models tested on the A2 validation set across both the R20P and C500 subsets. Figure 3 visualizes multiple samples generated by EP-Diffuser.
We observe a reversed ranking when evaluating prediction plausibility and . For example, on R20P, despite its smaller model size ( million parameters) and the largest (), EP-Diffuser achieves the highest realism meta score at and performs comparably to OptTrajDiff in terms of agent interaction and map adherence, while also maintaining high coverage (). Notably, EP-Diffuser demonstrates superior agent kinematics (), outperforming OptTrajDiff () and FMAE-MA ().
As an illustrative example, Figure 4 visualizes the vehicle kinematics during a left-turn maneuver in A2. EP-Diffuser produces the most plausible agent kinematics according to the measures in [38].
Figure 5 reports the realism meta scores and inference time of diffusion-based models over a range of DDIM denoising steps, with FMAE-MA as the regression-based baseline for comparison. Across all denoising step configurations, EP-Diffuser consistently outperforms OptTrajDiff in both metrics, achieving higher realism scores with lower inference time. While EP-Diffuser’s inference time exceeds FMAE-MA’s, it reaches peak performance with 5 denoising steps in just , making it viable for real-time applications.
@@ IV-C OoD Testing on WO
In the OoD Setting, the model trained on A2 is asked to generate traffic scene continuations for independent scenes taken from the WO dataset – thus removing any sort of shared bias between training and test data. The OoD testing results are presented in the “A2/WO” sections in Table II.
In the OoD setting, EP-Diffuser maintains the top scores in realism meta and agent kinematics. Additionally, it also demonstrates improved performance by achieving the best agent interactions and map adherence scores on both subsets. Furthermore, EP-Diffuser outperforms FMAE-MA in and closely matches OptTrajDiff ( vs. on C500, respectively) with high coverage (). These results demonstrate EP-Diffuser’s ability to learn robust, transferable representations from training data, highlighting its enhanced generalization beyond dataset-specific patterns.
IV-D Ablation Study
To evaluate the impact of polynomial representations, we introduce a variant, Seq-Diffuser, which replaces polynomial inputs and outputs in EP-Diffuser with sequence-based representations. Only minor modifications to the model architecture are required to accommodate the increased dimensionality of the sequence-based representation. The training and diffusion settings are identical for both EP-Diffuser and Seq-Diffuser. Comparative results are presented in Table II.
While Seq-Diffuser achieves better in-distribution accuracy with lower (e.g. vs. on R20P), it exhibits a notable degradation in both the plausibility and diversity of the generated traffic scenes across both subsets. Moreover, Seq-Diffuser generalizes poorly, as reflected by its lower performance in plausibility and higher OoD (e.g. vs. on R20P). It also incurs a higher inference time, as shown in Figure 5.
These results indicate that, although sequence-based representations may capture finer-grained details beneficial for pointwise accuracy, polynomial representations lead to better plausibility, diversity, and OoD generalization in multi-agent scene generation.